# gRNAde: Geometric Deep Learning for 3D RNA inverse design

### v6

@@ Abstract
Computational RNA design tasks are often posed as inverse problems, where sequences are designed based on adopting a single desired secondary structure without considering 3D geometry and conformational diversity. We introduce gRNAde, a geometric RNA design pipeline operating on 3D RNA backbones to design sequences that explicitly account for structure and dynamics. gRNAde uses a multi-state Graph Neural Network and autoregressive decoding to generates candidate RNA sequences conditioned on one or more 3D backbone structures where the identities of the bases are unknown. On a single-state fixed backbone re-design benchmark of 14 RNA structures from the PDB identified by Das et al. (2010), gRNAde obtains higher native sequence recovery rates (56% on average) compared to Rosetta (45% on average), taking under a second to produce designs compared to the reported hours for Rosetta. We further demonstrate the utility of gRNAde on a new benchmark of multi-state design for structurally flexible RNAs, as well as zero-shot ranking of mutational fitness landscapes in a retrospective analysis of a recent ribozyme. Open source code: github.com/chaitjo/geometric-rna-design
@@ 1 Introduction
Challenges of RNA modelling. Despite the promises of RNA therapeutics, proteins have instead been the primary focus in the 3D biomolecular modelling community. Availability of a large number of protein structures from the PDB combined with advances in deep learning for structured data (Bronstein et al., 2021; Duval et al., 2023) have revolutionized protein 3D structure prediction (Jumper et al., 2021) and rational design (Dauparas et al., 2022; Watson et al., 2023). Applications of deep learning for computational RNA design are underexplored compared to proteins due to paucity of 3D structural data (Schneider et al., 2023). Most tools for RNA design primarily focus on secondary structure without considering 3D geometry (Churkin et al., 2018) and use non-learnt algorithms for aligning 3D RNA fragments (Han et al., 2017; Yesselman et al., 2019), which can be restrictive due to the hand-crafted nature of the heuristics used.
In addition to limited 3D data for training deep learning models, the key technical challenge is that RNA is more dynamic than proteins. The same RNA can adopt multiple distinct conformational states to create and regulate complex biological functions (Ganser et al., 2019; Hoetzel & Suess, 2022; Ken et al., 2023). Computational RNA design pipelines must account for both the 3D geometric structure and conformational flexibility of RNA to engineer new biological functions.
Our contributions. This paper introduces gRNAde, a geometric deep learning-based pipeline for RNA inverse design conditioned on 3D structure, analogous to ProteinMPNN for proteins (Dauparas et al., 2022). As illustrated in Figure 1, gRNAde generates candidate RNA sequences conditioned on one or more backbone 3D conformations, enabling both single- and multi-state fixed-backbone sequence design. We demonstrate the utility of gRNAde for the following design scenarios:
  • Improved performance and speed over Rosetta. We compare gRNAde to Rosetta (Leman et al., 2020), the state-of-the-art physically based tool for 3D RNA inverse design, for single-state fixed backbone design of 14 RNA structures of interest from the PDB identified by Das et al. (2010). We obtain higher native sequence recovery rates with gRNAde (56% on average) compared to Rosetta (45% on average). Additionally, gRNAde is significantly faster than Rosetta for inference; e.g. sampling 100+ designs in 1 second for an RNA of 60 nucleotides on an A100 GPU (<10 seconds on CPU), compared to the reported hours for Rosetta on CPU.
@@ 2.1 RNA conformational ensembles as geometric multi-graphs
    Edge features for each edge from node $j$ to $i$ are initialized as: (a) the unit vector from the source to destination node, $\vec{{\bm{x}}}_{j}-\vec{{\bm{x}}}_{i}$; (b) the distance in 3D space, $\|\vec{{\bm{x}}}_{j}-\vec{{\bm{x}}}_{i}\|_{2}$, encoded by 32 radial basis functions; and (c) the distance along the backbone, $j-i$, encoded by 32 sinusoidal positional encodings.
@@
The input to gRNAde is thus a set of geometric graphs $\{\mathcal{G}^{(1)},\dots,\mathcal{G}^{(k)}\}$ which is merged into what we term a ‘multi-graph’ representation of the conformational ensemble, $\mathcal{M}=({\bm{A}},{\bm{S}},\vec{{\bm{V}}})$, by stacking the set of scalar features $\{{\bm{S}}^{(1)},\dots,{\bm{S}}^{(k)}\}$ into one tensor ${\bm{S}}\in\mathbb{R}^{n\times k\times f}$ along a new axis for the set size $k$. Similarly, the set of vector features $\{\vec{{\bm{V}}}^{(1)},\dots,\vec{{\bm{V}}}^{(k)}\}$ is stacked into one tensor $\vec{{\bm{V}}}\in\mathbb{R}^{n\times k\times f^{\prime}\times 3}$. Lastly, the set of adjacency matrices $\{{\bm{A}}^{(1)},\dots,{\bm{A}}^{(k)}\}$ are merged via a union $\cup$ into one single joint adjacency matrix ${\bm{A}}$.
@@ 2.2 Multi-state GNN for encoding conformational ensembles
A simple sum or average pooling does not introduce any new learnable parameters to the pipeline and is flexible to handle a variable number of conformations, enabling both single-state and multi-state design with the same model.
@@ 4.1 Single-state RNA design benchmark
We set out to compare gRNAde to Rosetta, a state-of-the-art physically based toolkit for biomolecular modelling and design (Leman et al., 2020). We reproduced the benchmark setup from Das et al. (2010) for Rosetta’s fixed backbone RNA sequence design workflow on 14 RNA structures of interest from the PDB, which mainly includes riboswitches, aptamers, and ribozymes (full listing in Table 2). We trained gRNAde on the single-state split detailed in Section 3, explicitly excluding the 14 RNAs as well as any structurally similar RNAs in order to ensure that we fairly evaluate gRNAde’s generalization abilities vs. Rosetta.
gRNAde improves sequence recovery over Rosetta. In Figure 4, we compare gRNAde’s native sequence recovery for single-state design with numbers taken from Das et al. (2010) for Rosetta, FARNA (a predecessor of Rosetta), ViennaRNA (the most popular 2D inverse folding method), and RDesign (Tan et al., 2023) (a concurrent GNN-based RNA inverse folding model). gRNAde has higher recovery of 56% on average compared to 45% for Rosetta, 32% for FARNA, 27% for ViennaRNA, and 43% for RDesign. See Appendix Table 2 for per-RNA results and Appendix B for full results on the single-state test set of 100 RNAs.
@@ 5 Conclusion
Limitations. Key avenues for future development of gRNAde include supporting multiple interacting chains, accounting for partner molecules with RNAs, and supporting negative design against undesired conformations. We discuss practical tradeoffs to using gRNAde in real-world RNA design scenarios in Appendix F, including limitations due to the current state of 3D RNA structure prediction tools.
@@ Acknowledgements
We would like to thank Roger Foo, Phillip Holliger, Alex Borodavka, Rhiju Das, Janusz Bujnicki, Edoardo Gianni, Ben Porebski, Michael Mohsen, Christian Choe, and John Boom for helpful comments and discussions. CKJ was supported by the A*STAR Singapore National Science Scholarship (PhD). SVM was supported by the UKRI Centre for Doctoral Training in Application of Artificial Intelligence to the study of Environmental Risks (EP/S022961/1). AM was supported by a U.S. NSF grant (DBI2308699) and two U.S. NIH grants (R01GM093123 and R01GM146340). This research was partially supported by Google TPU Research Cloud and Cambridge Dawn Supercomputer Pioneer Project compute grants.
@@ Appendix A Related Work
RNA inverse folding. Most tools for RNA inverse folding focus on secondary structure without considering 3D geometry (Churkin et al., 2018; Runge et al., 2019) and approach the problem from the lens of energy optimisation (Ward et al., 2023). Rosetta fixed backbone re-design (Das et al., 2010) is the only energy optimisation-based approach that accounts for 3D structure. Deep neural networks such as gRNAde can incorporate 3D structural constraints and are orders of magnitude faster than optimisation-based approaches; this is particularly attractive for high-throughput design pipelines as solving the inverse folding optimisation problem is NP hard (Bonnet et al., 2020).
@@ Appendix C Additional Results
Table 2: Full results for Figure 4 comparing gRNAde to Rosetta, FARNA, ViennaRNA and RDesign for single-state design on 14 RNA structures of interest identified by Das et al. (2010). Rosetta and FARNA recovery values are taken from Das et al. (2010), Supplementary Table 2.
@@
Appendix D 3D Visualisation of gRNAde Designs
(a) self-consistency TM-score
(b) self-consistency GDT_TS
Figure 8: 3D self-consistency scores for 3 representative RNAs from Das et al. (2010). We use RhoFold to ‘forward fold’ 100 designs sampled at temperature = 0.5 and plot self-consistency TM-score and GDT_TS. Each dot corresponds to one designed sequence and is coloured by gRNAde’s perplexity (normalised per RNA). Designs with lower relative perplexity generally have higher 3D self-consistency and can be considered more ‘designable’. Dotted lines represent TM-score and GDT_TS thresholds of 0.45 and 0.50, repsectively. Pairs of structures scoring higher than the threshold correspond to roughly the same fold.
Design 1:
GGCAAGUAAUCCCUACGCUAUG
GGUAGGGAGUCUCAGCAGUGAC
CCGUAAAGUUACUACCUUGCCC
perplexity: 1.3097
recovery: 0.5909 (27 edits)
sc2D = 0.9227
scRMSD = 1.3839
scTM = 0.8309
scGDT = 0.8295
Design 2:
CGGUGGUAAGCCCAACGCUAGG
GGUUGGGCGUCUCAGCACAGUC
CCGUAAAGAUUGUACCCACCGG
perplexity: 1.3815
recovery: 0.4091 (37 edits)
sc2D = 0.9227
scRMSD = 2.1249
scTM = 0.6874
scGDT = 0.6780
Design 3:
AGCAAGUAAUGCCAUCGCUAUG
GGAUGGUAGUGUCAGCACUGAC
CCUUAAAGUUAGUACCUUGCUU
perplexity: 1.4247
recovery: 0.5152 (30 edits)
sc2D = 0.9227
scRMSD = 3.2131
scTM = 0.5118
scGDT = 0.5265
Figure 9: Cherry-picked designs for Guanine riboswitch aptamer (PDB: 4FE5, sequence: GGACAUAUAAUCGCGUGGAUAUGGCACGCAAGUUUCUACCGGGCACCGUAAAUGUCCGACUAUGUCC). We show the RhoFold-predicted 3D structure in colour overlaid on the groundtruth structure in grey. Designs recover the base pairing patterns and tertiary structure of the RNA, as measured by high self-consistency score. gRNAde’s perplexity is correlated well with 3D self-consistency scores and can be useful for ranking designs.
Design 1:
GGGGCUCCGGCGACGCAGUCGA
AAGCCCAGCAGUACCAAGCCUC
AGGGGAAACUUUGAGGUGGCCU
AACAAAGGAUACGGUAAUAAGC
UGCGGGAAAAGGUUGUAAGCCG
GAGCGAAGACCUAAGGCACCGC
UUUUGGCGGUGCUAUGGUUGAA
GUUAA
perplexity: 1.2462
recovery: 0.7170 (44 edits)
sc2D = 0.8301
scRMSD = 5.4562
scTM = 0.6481
scGDT = 0.4465
Design 2:
GGGGUACCGGCGACGCAGUCG
AAUGCCCUGUGGUACCAAGCC
CCGGGGGAAACUUCGGGGUGG
CCUUACCAAGGACACGGUAAU
AAGCCACGGGAAAUGGUUGUA
AGCCGGUCCGAAGCCCUAAGG
CCGCGCUUUGGGCGCGGCUAU
GGGUGAAGGCAA
perplexity: 1.3273
recovery: 0.6226 (58 edits)
sc2D = 0.6896
scRMSD = 6.7239
scTM = 0.6300
scGDT = 0.4513
Design 3:
GAGGCCACGGCAACGCAGUC
UAACGCCCUGUGGUACCAAG
UCUUAGGAGAAAUUUUAAGA
UGGCCUAAUAAAGGAUAUGG
UAAUAAGCCACGGGAAAAGG
UUGUAAGACGUGACGAAGUC
CUAAGGCCACAGUUUUGCUG
UGGCUAUGGAUGGAGUACA
perplexity: 1.3204
recovery: 0.7044 (45 edits)
sc2D = 0.7922
scRMSD = 8.8211
scTM = 0.4582
scGDT = 0.2909
Figure 10: Cherry-picked designs for Tetrahymena Ribozyme P4-P6 domain (PDB: 2R8S, sequence: GGAAUUGCGGGAAAGGGGUCAACAGCCGUUCAGUACCAAGUCUCAGGGGAAACUUUGAGAUGGC
CUUGCAAAGGGUAUGGUAAUAAGCUGACGGACAUGGUCCUAACACGCAGCCAAGUCCUAAGUCAACAGAUC
UUCUGUUGAUAUGGAUGCAGUUCA
Design 1:
GUCAAACGCAGCCGAAA
GCGCGAUAGUCCCAGGAA
perplexity = 1.6237
recovery = 0.4571 (16 edits)
sc2D = -0.0074
scRMSD = 3.9505
scTM = 0.2597
scGDT = 0.4786
Design 2:
GGCAAACGCGGCCGAAA
GCGCGUGAGUCCCCGGAC
perplexity = 1.6630
recovery = 0.4857 (16 edits)
sc2D = -0.0099
scRMSD = 3.3549
scTM = 0.2526
scGDT = 0.5000
Design 3:
CGUAGUCGGAGCCGAAG
GGCCGUUAGUCCCAGGAG
perplexity = 1.7020
recovery = 0.4000 (17 edits)
sc2D = 0.4035
scRMSD = 16.4102
scTM = 0.0319
scGDT = 0.0571
Figure 11: Cherry-picked designs for Vitamin B12 binding aptamer (PDB: 1ET4, sequence: GGAACCGGUGCGCAUAACCACCUCAGUGCGAGCAA).
Appendix E Additional Figures

### v7

@@ Abstract
Computational RNA design tasks are often posed as inverse problems, where sequences are designed based on adopting a single desired secondary structure without considering 3D conformational diversity. We introduce gRNAde, a geometric RNA design pipeline operating on 3D RNA backbones to design sequences that explicitly account for structure and dynamics. gRNAde uses a multi-state Graph Neural Network and autoregressive decoding to generates candidate RNA sequences conditioned on one or more 3D backbone structures where the identities of the bases are unknown. On a single-state fixed backbone re-design benchmark of 14 RNA structures from the PDB identified by Das et al. (2010), gRNAde obtains higher native sequence recovery rates (56% on average) compared to Rosetta (45% on average), taking under a second to produce designs compared to the reported hours for Rosetta. We further demonstrate the utility of gRNAde on a new benchmark of multi-state design for structurally flexible RNAs, as well as zero-shot ranking of mutational fitness landscapes in a retrospective analysis of a recent ribozyme. Experimental wet lab validation on 10 different structured RNA backbones finds that gRNAde has a success rate of 50% at designing pseudoknotted RNA structures, a significant advance over 35% for Rosetta. Open source code and tutorials are available at: github.com/chaitjo/geometric-rna-design
@@ 1 Introduction
Challenges of RNA modelling. Despite the promises of RNA therapeutics, proteins have been the primary focus in the 3D biomolecular modelling community. Availability of a large number of protein structures from the PDB combined with advances in deep learning for structured data (Bronstein et al., 2021; Duval et al., 2023) have revolutionized protein 3D structure prediction (Jumper et al., 2021) and rational design (Dauparas et al., 2022; Watson et al., 2023). Applications of deep learning for RNA design are underexplored compared to proteins due to paucity of 3D structural data (Schneider et al., 2023). Most tools for RNA design primarily focus on secondary structure without considering 3D geometry (Churkin et al., 2018) or use non-learnt algorithms for aligning 3D RNA fragments (Han et al., 2017; Yesselman et al., 2019), which can be restrictive due to the hand-crafted nature of the heuristics used. In addition to limited 3D data, the key technical challenge is that RNA is generally more dynamic than proteins. The same RNA can adopt multiple distinct conformational states to create and regulate complex biological functions (Ganser et al., 2019; Hoetzel & Suess, 2022; Ken et al., 2023). Computational RNA design pipelines must account for both the 3D geometric structure and conformational flexibility of RNA to engineer new biological functions.
Our contributions. This paper introduces gRNAde, a geometric deep learning-based pipeline for RNA inverse design. As illustrated in Figure 1, gRNAde generates candidate RNA sequences conditioned on one or more 3D backbone structures, enabling both single- and multi-state fixed-backbone sequence design. We demonstrate the utility of gRNAde for the following design scenarios:
  • Improved performance and speed over Rosetta. We compare gRNAde to Rosetta (Leman et al., 2020), the state-of-the-art physics-based tool for 3D RNA inverse design, for single-state fixed backbone design of 14 RNA structures of interest from the PDB identified by Das et al. (2010). We obtain higher native sequence recovery rates with gRNAde (56% on average) compared to Rosetta (45% on average). Additionally, gRNAde is significantly faster than Rosetta for inference; e.g. sampling 100+ designs in 1 second for an RNA of 60 nucleotides on an A100 GPU (<10 seconds on CPU), compared to the reported hours for Rosetta on CPU.
@@
  • Wet lab validated. As part of Eterna’s OpenKnot Round 6, 200 gRNAde-designed RNAs for diverse backbones were independently validated in a wet lab via SHAPE chemical mapping experiments. gRNAde demonstrated an overall success rate of 50% at designing RNAs with desired pseudoknotted structures, which is a significant improvement over Rosetta with 35%.
@@ 2.1 RNA conformational ensembles as geometric multi-graphs
    Edge features from node $j$ to $i$ are initialized as: (a) the unit vector from the source to destination node, $\vec{{\bm{x}}}_{j}-\vec{{\bm{x}}}_{i}$; (b) the distance in 3D space, $\|\vec{{\bm{x}}}_{j}-\vec{{\bm{x}}}_{i}\|_{2}$, encoded by 32 radial basis functions; and (c) the distance along the backbone, $j-i$, encoded by 32 sinusoidal positional encodings.
@@
The input to gRNAde is thus a set of geometric graphs $\{\mathcal{G}^{(1)},\dots,\mathcal{G}^{(k)}\}$ which is merged into a ‘multi-graph’ representation of the conformations, $\mathcal{M}=({\bm{A}},{\bm{S}},\vec{{\bm{V}}})$, by stacking the set of scalar features $\{{\bm{S}}^{(1)},\dots,{\bm{S}}^{(k)}\}$ into one tensor ${\bm{S}}\in\mathbb{R}^{n\times k\times f}$ along a new axis for the set size $k$. Similarly, the set of vector features $\{\vec{{\bm{V}}}^{(1)},\dots,\vec{{\bm{V}}}^{(k)}\}$ is stacked into one tensor $\vec{{\bm{V}}}\in\mathbb{R}^{n\times k\times f^{\prime}\times 3}$. Lastly, the set of adjacency matrices $\{{\bm{A}}^{(1)},\dots,{\bm{A}}^{(k)}\}$ are merged via union $\cup$ into a joint adjacency matrix ${\bm{A}}$.
@@ 2.2 Multi-state GNN for encoding conformational ensembles
A simple sum or average pooling does not introduce any new learnable parameters to the pipeline and is flexible to handle a variable number of conformations, enabling both single-state and multi-state design with the same model. In Appendix B, we also explore more expressive geometric set pooling functions (Maron et al., 2020).
@@ 4.1 Single-state RNA design benchmark
We set out to compare gRNAde to Rosetta, a state-of-the-art physics-based toolkit for biomolecular modelling and design (Leman et al., 2020). We reproduced the benchmark setup from Das et al. (2010) for Rosetta’s fixed backbone RNA sequence design workflow on 14 RNA structures of interest from the PDB, which mainly includes riboswitches, aptamers, and ribozymes (full listing in Table 2). We trained gRNAde on the single-state split detailed in Section 3, explicitly excluding the 14 RNAs as well as any structurally similar RNAs in order to ensure that we fairly evaluate gRNAde’s generalization abilities vs. Rosetta.
gRNAde improves sequence recovery over Rosetta. In Figure 4, we compare gRNAde’s native sequence recovery for single-state design with numbers taken from Das et al. (2010) for Rosetta, FARNA (a predecessor of Rosetta), ViennaRNA (the most popular 2D inverse folding method), and RDesign (Tan et al., 2023). RDesign is a concurrent deep learning-based 3D inverse folding model which uses invariant GNN layers and non-autoregressive decoding without sampling; see Appendix F for details. gRNAde has higher recovery of 56% on average compared to 45% for Rosetta, 32% for FARNA, 27% for ViennaRNA, and 43% for RDesign.
@@ 4.3 Zero-shot ranking of RNA fitness landscape
Figure 7: Wet lab validation of gRNAde designs in Eterna’s OpenKnot Round 6. Each point represents the OpenKnot score (higher is better) for a designed RNA sequence given one of 10 target RNA backbones (10 designs per target per method). Wildtype sequences were included as sanity checks and upper bounds on performance. 50% of gRNAde designs have a score above 80, which is the threshold for a successful design. The median gRNAde design obtains a score of 80, while Rosetta designs have a median score of 76, with a 35% sucess rate. See Appendix Figure 10 for per-puzzle results. Note that we have obtained permission from Eterna organizers to share this data.
5 Independent wet lab validation of gRNAde designs
Finally, we present the results of independent wet lab validation of gRNAde via Eterna, an online platform for computational RNA design. Eterna regularly releases new RNA design tasks to a global community of researchers and citizen-scientists who design sequences using expert intuition or computational tools (such as gRNAde and Rosetta). The designs are then experimentally validated at Stanford University via high-throughput SHAPE assays which measures the reactivity of an RNA sequence to a chemical modifier, as described in He et al. (2024).
Eterna OpenKnot Round 6. As part of OpenKnot Round 6, we submitted gRNAde designs for 10 target RNA backbones: SARS-CoV-2 frame shift element, Tetrahydrofolate riboswitch, GMP-II riboswitch, SAM riboswitch, HCV internal ribosome entry site, synthetic kissing loop structure, donggang dumbell, telomerase ribozyme, HDV ribozyme, and CPEB3 ribozyme. We submitted a total of 20 designs for each backbone via two approaches: (1) 10 partial designs, where parts of the wildtype sequence are kept fixed; and (2) 10 full designs, where the entire sequence is designed. In addition to designs from participants including gRNAde, the organizers also evaluated the wildtype sequence for each RNA as a sanity check for their assay, as well as two variants of Rosetta’s RNA inverse folding protocol (Das et al., 2010).
OpenKnot score determines success. For each design, the organizers compute an OpenKnot Score (between 0–100) to measure how likely the sequence is to form the target 3D structure. A score above 80 is estimated to be highly likely to form the target structure (scores below 80 may also be successful, but their SHAPE reactivity profiles cannot determine this). The score was developed to aid the discovery and characterisation of pseudoknotted structural elements in RNAs using SHAPE data as part of the OpenKnot challenge series.
gRNAde has 50% success rate and outperforms Rosetta. In Figure 7, we plot the distribution of OpenKnot scores across all target RNAs for gRNAde and Rosetta designs as well as the wildtype control sequences. 52% of gRNAde partial designs and 49% of gRNAde full designs obtained OpenKnot scores greater than 80, compared to 35% for the best performing Rosetta variant. The median gRNAde design has a score of 80, while the median Rosetta design scores 76. See Appendix Figure 10 for per-puzzle results where we find that gRNAde designs obtain higher scores than wildtype sequences for 3 targets, suggesting that designed sequences can have higher likelihood of forming the target structure compared to sequences observed in nature.
@@ 6 Conclusion
Limitations. Key avenues for future development of gRNAde include supporting multiple interacting chains, accounting for partner molecules with RNAs, and supporting negative design against undesired conformations. We discuss practical tradeoffs to using gRNAde in real-world RNA design scenarios in Appendix E, including limitations due to the current state of 3D RNA structure prediction tools. Finally, we are hopeful that advances in RNA structure determination and computationally assisted cryo-EM (Kappel et al., 2020; Bonilla & Kieft, 2022) will further increase the amount of RNA structures available for training geometric deep learning models in the future.
@@ Acknowledgements
We would like to thank Roger Foo, Phillip Holliger, Alex Borodavka, Rhiju Das, Janusz Bujnicki, Edoardo Gianni, Ben Porebski, Samantha Kwok, Michael Mohsen, Christian Choe, and John Boom for helpful comments and discussions. CKJ was supported by the A*STAR Singapore National Science Scholarship (PhD) and Qualcomm Innovation Fellowship. SVM was supported by the UKRI Centre for Doctoral Training in Application of Artificial Intelligence to the study of Environmental Risks (EP/S022961/1). AM was supported by a U.S. NSF grant (DBI2308699) and two U.S. NIH grants (R01GM093123 and R01GM146340). This research was partially supported by Google TPU Research Cloud and Cambridge Dawn Supercomputer Pioneer Project compute grants.
@@ Appendix A Related Work
RNA inverse folding. Most tools for RNA inverse folding focus on secondary structure without considering 3D geometry (Churkin et al., 2018; Runge et al., 2019; Cole et al., 2024) and approach the problem from the lens of energy optimisation (Ward et al., 2023). Rosetta fixed backbone re-design (Das et al., 2010; Leman et al., 2020) is the only energy optimisation-based approach that accounts for 3D structure. Rosetta aims to find the lowest energy RNA sequence for a given structure: (1) Given a starting point such as a target backbone and a random/heuristic starting sequence, Rosetta uses Markov Chain Monte Carlo methods to sample ‘moves’ which are random changes to the sequence (whether a base is an A, U, G, or C) or side chain rotamer (the orientation of the base). (2) Next, Rosetta updates the all-atom 3D structure of the RNA backbone and uses an all-atom energy function to accept or reject a move based on whether it stabilises the target backbone. This process is repeated until convergence with simulated annealing.
Deep neural networks such as gRNAde can incorporate 3D structural constraints and produce designs in a single forward pass with GPU acceleration, which is orders of magnitude faster than optimisation-based approaches. This is particularly attractive for high-throughput design pipelines as solving the inverse folding optimisation problem is NP hard (Bonnet et al., 2020). Independent of our work, Tan et al. (2023) also developed a contemporaneous deep learning-based 3D RNA inverse folding model.
@@ Appendix B Ablation Study
In order to study the expressivity of pooling in the multi-state setting, we ablate the set pooling function used in the multi-state GNN: Deep Set pooling (DS) as well as the more expressive Deep Symetric Set pooling (DSS, Maron et al. (2020)). DS pooling is described in Equation 3. In DSS pooling, after each encoder GNN layer, we (1) aggregate node embeddings from each single state representation into a pooled multi-state representation per node, (2) apply a Geometric Vector Perceptron update on the multi-state representation, and (3) add the updated multi-state representation back to each single state node representation. DSS pooling marginally improves performance on out-of-distribution test sets for both the single- and multi-state splits. We notice that DSS models fit the training data significantly better (final training loss goes from 0.40 to 0.36 for 3 states). While DSS pooling is more expressive that DS pooling, it also adds 200K more parameters to the model and doubles training iteration time (4 mins to 8 mins for 3 states).
@@ Appendix C Additional Results
Table 2: Full results for Figure 4 comparing gRNAde to Rosetta, FARNA and ViennaRNA for single-state design on 14 RNA structures of interest identified by Das et al. (2010). Rosetta and FARNA recovery values are taken from Das et al. (2010), Supplementary Table 2.
@@
Figure 8: gRNAde compared to RDesign for single-state design.
@@
Figure 10: Per-input wet lab validation results of gRNAde designs. Notably, for the HDV ribozyme, CPEB3 ribozyme and donggang dumbell backbone structures, gRNAde designs tend to obtain higher OpenKnot scores than the wildtype sequences, which suggests that designed sequences can have higher likelihood of forming the pseudoknotted structure compared to naturally observed sequences.
Appendix D Additional Figures